# Lightening the Load: A Cluster-Based Framework for A Lower-Overhead, Provable Website Fingerprinting Defense

### v1

@@ B-C Cluster Diversity Evaluation
As shown in the figure, the empirical purity values closely track the $1/k$ curve, indicating that our clustering procedure approaches the theoretical optimum in terms of diversity. This result confirms that the Tamaraw-specific distance function naturally leads to anonymity sets composed of traffic patterns from multiple websites, supporting the desired $l$-diversity property of the defense.
Appendix C Proofs
C-A Security Bound for Adaptive Tamaraw
@@
We use two padding operators:
  • $D_{\text{glob}}(\,\cdot\,;\,p_{\mathrm{in}}^{g},p_{\mathrm{out}}^{g})$ applies the global constant–rate schedule with parameters $p_{\mathrm{in}}^{g},p_{\mathrm{out}}^{g}$.
  • $D_{S_{i}}(\,\cdot\,;\,p_{\mathrm{in}}^{(S_{i})},p_{\mathrm{out}}^{(S_{i})},L)$ applies the AS–specific schedule chosen for $S_{i}$.
Given a raw trace $f\in S_{i}$ the full defense acts in two stages:
$\widehat{f}=D(f)=\begin{cases}D_{\text{glob}}\!\bigl{(}f(t);\;p_{\mathrm{in}}^{g},p_{\mathrm{out}}^{g}\bigr{)},&0\leq t<t_{S_{i}},\\[6.0pt] D_{S_{i}}\!\bigl{(}f(t);\;p_{\mathrm{in}}^{(S_{i})},p_{\mathrm{out}}^{(S_{i})},L\bigr{)},&t\geq t_{S_{i}}.\end{cases}$
We write this concisely as a labelled transition
$D_{\text{glob}}\;\xrightarrow[\;t_{S_{i}}\;]{}D_{S_{i}}.$
@@
Lemma C.1 (Post-switch non-uniform weighted $\delta$).
@@
Theorem C.2 (Global non-uniformly weighted $\delta$–non-injectivity).
@@
Recall from Lemma C.1 that for each anonymity set $S_{i}$, the attacker’s success rate over its defended outputs is bounded by $\bar{A}(S_{i})=\mathbb{E}_{f^{\prime}\sim S_{i}}\left[\frac{1}{\tilde{\delta}(f^{\prime})}\right]$, and thus the mapping $D_{S_{i}}$ is non-uniformly $\delta_{i}$-weighted-non-injective, where $\delta_{i}=1/\bar{A}(S_{i})$.
@@
Appendix D Modified CAST Algorithm
This section presents the algorithmic components used in Adaptive Tamaraw. In particular, we provide the full pseudocode for our modified version of the CAST algorithm, which plays a central role in obtaining the intra-website patterns, mentioned in Section V-C.
Algorithm 2 Intra-Website Pattern Detection via Modified CAST
1:A set of traffic traces $\{F_{1},F_{2},\ldots,F_{n}\}$ from a single website
2:A set of balanced clusters $\mathcal{P}=\{\mathcal{C}_{1},\mathcal{C}_{2},\ldots,\mathcal{C}_{m}\}$
3:Step 1: Similarity Matrix Construction
4:for each trace $F_{i}$ do
5:  Compute Euclidean distances $d(F_{i},F_{j})$ for all $j$
6:  Determine local scale $\sigma_{i}$ as the distance from $F_{i}$ to its $K$th nearest neighbor (with $K=7$)
7:end for
8:Compute the similarity matrix $S$, where:
[…]
10:Step 2: Initial Clustering with Modified CAST
11:Initialize all traces as unassigned.
12:while there exist unassigned traces do
13:  for each unassigned trace $F_{x}$ do
14:   for each existing cluster $\mathcal{C}$ do
15:     Compute affinity:
16:     $a_{\mathcal{C}}(F_{x})=\frac{1}{|\mathcal{C}|}\sum_{F_{y}\in\mathcal{C}}S(F_{x},F_{y})$
17:   end for
18:   Assign $F_{x}$ to the cluster $\mathcal{C}^{*}$ for which $a_{\mathcal{C}^{*}}(F_{x})$ is maximum, provided that
19:   $a_{\mathcal{C}^{*}}(F_{x})\geq\tau$, where $\tau$ is a dynamically chosen affinity threshold.
20:   if no existing cluster meets the threshold then
21:     Create a new cluster with $F_{x}$ as the seed.
22:   end if
23:  end for
24:  Perform a cleaning step: For every trace, reassign it to the cluster with which it has the highest affinity if different from its current cluster.
25:end while
26:Step 3: Post-Processing
27:while the total number of clusters $\text{MaxClusters}$ do
28:  Sort clusters in $\mathcal{P}$ by size.
29:  Merge the smallest cluster into the adjacent cluster that minimizes the expansion ratio, defined as
30:  $\phi(\mathcal{C})=\frac{\mathrm{cut}(\mathcal{C})}{\mathrm{vol}(\mathcal{C})}$,
31:  where for a cluster $\mathcal{C}$,
32:  $\mathrm{cut}(\mathcal{C})=\sum_{F\in\mathcal{C}}\sum_{F^{\prime}\notin\mathcal{C}}d(F,F^{\prime})$,
33:  and
34:  $\mathrm{vol}(\mathcal{C})=\sum_{F,F^{\prime}\in\mathcal{C}}d(F,F^{\prime})$.
35:  Update the clustering accordingly.
36:end while
37:return the final set of clusters $\mathcal{P}$.

### v2

@@ C-A Cluster Diversity Evaluation
C-B CAST Modifications
To assess the effect of our CAST algorithm changes, we performed an ablation study measuring each modification’s impact on clustering metrics using the Sirinam et al. dataset (Table VII). The baseline CAST creates highly fragmented clusters, with over 81% containing fewer than 20 traces. Adding a dynamic threshold increases average cluster size, while cleaning and post-processing improve size balance. Cleaning reduces small clusters from 64.3% to 13.6%, and post-processing merges remaining ones to 0.2%, yielding a final size ratio of 3.83 and an average cluster size of 200. Together, these updates produce stable, well-balanced traffic pattern clusters vital to our defense framework.
TABLE VII: Effects of CAST modifications on clustering quality.
Variant Average Small Cluster Largest/
Cluster Size % Smallest
Baseline (Raw) 13 81.9 200
+ Dynamic Threshold 70 64.3 633
+ Cleaning Step 74 13.6 27.7
+ Post Processing 200 0.2 3.83
C-C Sensitivity Analysis on $\alpha$
The early-switching confidence threshold $\alpha$ controls the balance between efficiency and security in our adaptive defense by defining the safe timestamp $\tau_{s}$ where the system switches from global to local configuration. We derived anonymity sets from the training set and ran a sensitivity study on validation data. Lower $\alpha$ values trigger earlier, less confident switching, extending the use of efficient local parameters but increasing vulnerability, as predictions align less with true anomaly structures. As Table VIII shows, reducing $\alpha$ from 1.0 to 0.6 boosts overhead savings from $\sim$15% to $\sim$34%, while the $u_{s} atta$cker accuracy (introduced in Appendix B-B) rises from 34% to 41%. We chose $\alpha=0.9$ for its balanced trade-off.
TABLE VIII: Performance metrics across different alpha values for adaptive defense mechanism.
Metrics $\alpha$ 0.6 0.7 0.8 0.9 1.0
Overhead Improvement -34% -32% -27% -25% -15%
$\tau_{s}$ Aware Attacker Acc. 41% 39% 38% 36% 34%
C-D Webpage vs Pattern Analysis
To quantify the end-to-end advantage of clustering fine-grained traffic patterns rather than the conventional approach of clustering at the website level, we conducted an ablation study comparing the performance of Adaptive Tamaraw under both configurations, using a bucket length of $L=100$ on the Sirinam et al. dataset. Table IX presents the resulting average total overhead (bandwidth + time), confirming that operating at the pattern level consistently yields lower overhead across all tested values of $k$. For instance, at $k=2$, our fine-grained approach reduces overhead by 11 percentage points compared to the website-level baseline. This performance gain stems from the increased homogeneity of pattern-level anonymity sets.
TABLE IX: Combined Overheads across different granularity levels.
Granularity $k=2$ $k=7$ $k=30$
Website Level 219% 229% 243%
Pattern Level 205% 217% 235%
Appendix D Computational and Memory Overhead
We report the computational and memory costs of Adaptive Tamaraw, measured on a system with an AMD EPYC 9454 CPU and an NVIDIA H100 GPU (20 GB MIG). The main Holmes model is 8.21 MB, and each lightweight kFP model is 0.41 MB. Inference latency per decision point is 1.84 ms: 0.15 ms for site prediction and 1.69 ms for traffic pattern identification. Our ECDIRE procedure retains 4.09 safe timestamps per site, requiring about 160 MB total storage for 95 sites. As summarized in Table X, the defense is lightweight and suitable for real-time deployment.
TABLE X: Summary of computational and resource costs.
Metric Value
Holmes Model Size 8.21 MB
kFP Model Size (per instance) 0.41 MB
Total Inference Latency (per decision) 1.84 ms
Avg. Safe Timestamps per Site 4.09
Appendix E Proof of Security Bound for Adaptive Tamaraw
@@
Lemma E.1 (Post-switch non-uniform weighted $\delta$).
@@
Theorem E.2 (Global non-uniformly weighted $\delta$–non-injectivity).
@@
Recall from Lemma E.1 that for each anonymity set $S_{i}$, the attacker’s success rate over its defended outputs is bounded by $\bar{A}(S_{i})=\mathbb{E}_{f^{\prime}\sim S_{i}}\left[\frac{1}{\tilde{\delta}(f^{\prime})}\right]$, and thus the mapping $D_{S_{i}}$ is non-uniformly $\delta_{i}$-weighted-non-injective, where $\delta_{i}=1/\bar{A}(S_{i})$.